OpenStudy (anonymous):

help! question for you guys :) 6. Calculate the mass of 10ml of water in the graduated cylinder using 14.54 for volume and the density of water of 1.00 g/mL. Calculate the mass of the water using the formula for density: Density = mass volume

jimthompson5910 (jim_thompson5910):

so the volume of the water in the cylinder is 10 mL or is it 14.54 mL ?

OpenStudy (anonymous):

volume is 14.54

jimthompson5910 (jim_thompson5910):

ok so Density = Mass/Volume D = m/v 1 = m/14.54 1*14.54 = m 14.54 = m m = 14.54

jimthompson5910 (jim_thompson5910):

so the mass is 14.54 grams

OpenStudy (anonymous):

oh brilliant thanks so much for explaining that!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

quick question, if the density of the water is 1g/mL would i be multiplying 14.54 times 10 and not 1??

jimthompson5910 (jim_thompson5910):

mass = (Density)*(Volume) mass = (1 g/mL)*(14.54 mL) mass = 14.54 grams

jimthompson5910 (jim_thompson5910):

notice how the mL units match and cancel out

jimthompson5910 (jim_thompson5910):

this is of course assuming the given volume is 14.54 mL and not 10 mL

OpenStudy (anonymous):

ya i am using the volume from the previous problem (as instruscted to). . . just curious as to why the 1g/mL was being mentioned. . .

jimthompson5910 (jim_thompson5910):

that's the density of water basically it tells you how packed water is

jimthompson5910 (jim_thompson5910):

or it tells you how much mass is in one unit of volume

OpenStudy (anonymous):

ohhh ok gotcha :)